# MLP Memory: A Retriever-Pretrained Memory for Large Language Models

### v2

@@ Abstract
Modern approaches to enhancing Large Language Models’ factual accuracy and knowledge utilization face a fundamental trade-off: non-parametric retrieval-augmented generation (RAG) provides flexible access to external knowledge but suffers from high inference latency and shallow integration, while parametric fine-tuning methods like LoRA risk catastrophic forgetting and degraded general capabilities. In this work, we propose MLP Memory, a lightweight parametric module that learns to internalize retrieval patterns without explicit document access. By pretraining an MLP to imitate a NN retriever’s behavior on the entire pretraining dataset, we create a differentiable memory component that captures the benefits of retrieval-based knowledge access in a fully parametric form. Our architecture integrates this pretrained MLP Memory with Transformer decoders through simple probability interpolation, achieving 12.3% relative improvement on five question-answering benchmarks and 5.2 points absolute gain across nine general NLP tasks, while reducing hallucinations by up to 10 points on HaluEval. Moreover, MLP Memory delivers 2.5 faster inference than RAG with superior accuracy. Our findings show that learning retrieval patterns parametrically bridges the gap between efficient inference and effective knowledge access, offering a practical alternative to both RAG and fine-tuning approaches. 111Code and checkpoints will be released upon acceptance.
@@ 1 Introduction
Decoder-only architectures such as GPT (Brown et al., 2020), LLaMA (Grattafiori et al., 2024), Qwen (Qwen et al., 2025), and DeepSeek (Liu et al., 2024) have achieved remarkable success in various tasks, including open-ended text generation (OpenAI et al., 2024), code completion (Chen et al., 2021), image synthesis (Chen et al., 2020), and multimodal reasoning (Liu et al., 2023). However, despite their impressive capabilities, these models often struggle with effective knowledge utilization, producing responses that may be fluent but fail to accurately leverage the factual information encoded in their parameters.
@@
Current approaches to enhance knowledge utilization in LLMs face significant trade-offs. Retrieval-augmented generation (RAG) methods (Lewis et al., 2021; Peng et al., 2023; Gao et al., 2022; Izacard et al., 2022) dynamically fetch relevant documents to ground model outputs, providing flexible access to external knowledge sources. However, these non-parametric approaches introduce substantial inference latency through expensive nearest-neighbor searches and longer context from retrieved documents. They also suffer from shallow integration with the base model, as the retrieval component remains isolated from the LLM’s computational graph. Conversely, parametric adaptation methods such as continued pre-training (CPT) and LoRA (Hu et al., 2022) directly modify model weights to incorporate domain-specific knowledge. While computationally efficient at inference time, these approaches risk catastrophic forgetting of previously learned capabilities and often degrade performance on general tasks, requiring careful task-specific tuning that limits their broader applicability. Figure 2 illustrates how our approach differs fundamentally from both non-parametric retrieval methods and parametric adaptation approaches.
@@
In contrast to decoder-only LLMs, neuroscience research reveals a lateralized human brain where language processing is dominated by the left hemisphere while memory formation occurs in the hippocampus (Gazzaniga, 2005b; a; Douglas, 1967). This insight has inspired memory-augmented models in machine learning. Early approaches like Memory Networks (Weston et al., 2015) enabled read/write operations on external memory, while Sparse Access Memory introduced differentiable memory access schemes. However, these were task-specific with limited general applicability. In the LLM era, methods such as Memory Transformers (Burtsev et al., 2021) incorporate trainable memory tokens for global context, while AutoCompressors (Chevalier et al., 2023) compress long contexts into summary vectors. Nevertheless, these memory tokens primarily function as working memory supplements for context extension rather than long-term memory capable of retaining information from the entire training corpus.
@@
In this work, we propose an external memory for LLM that is pretrained to mimic a retriever on the entire pretraining dataset. Specifically, following the RAG setting in NN-LM (Khandelwal et al., 2020), this memory learns to map the LLM hidden state at a certain step to a vocabulary distribution matching the output of the NN retriever. During inference, the LLM’s native output is interpolated with the retriever-pretrained output from the external memory. Our resulting architecture, illustrated in Figure 4, consists of a transformer decoder and an external MLP memory, each pretrained separately with different pretraining tasks. For our pretrained external memory, we aim to achieve the following features simultaneously:
@@
Long-term memory, covering the whole pretraining corpus. While existing memory tokens serve primarily as working memory by storing local context for immediate use, our MLP memory functions as a long-term repository of generalizable knowledge acquired during the pretraining phase.
Experimental results demonstrate that MLP Memory significantly outperforms existing approaches across multiple dimensions. It achieves average relative improvements of 12.3% (Mistral-7B) and 7.8% (Llama2-7B) on five QA benchmarks, with WebQA showing exceptional gains (37.45% vs. 29.28% baseline). On nine general NLP tasks, it delivers a 5.2 points absolute improvement. MLP Memory also substantially reduces hallucinations on HaluEval, with accuracy improvements of 9.68, 10.08, and 2.14 points on dialogue, QA, and summarization tasks respectively. Most notably, it achieves 2.5 faster time-to-first-token than RAG and 5.6 faster than NN-LM, while maintaining constant inference speed regardless of corpus size, unlike retrieval methods whose latency scales with data size. Figure 1 illustrates MLP Memory’s performance gains and inference efficiency over baselines and Figure 3 demonstrates a case where MLP Memory correctly answers factual questions while RAG fails despite retrieving correct information. These results confirm that parametric compression of retrieval patterns offers a more efficient and effective alternative to explicit retrieval.
@@ 2 Preliminary: -nearest neighbors language model
The NN-LM (Khandelwal et al., 2020) augments a pre-trained LM by interpolating its parametric distribution with a non-parametric distribution from nearest neighbor retrieval. Given context , and denotes the next token. The next-token probability is:
@@
While NN-LM improves predictions through explicit memory, it suffers from substantial storage requirements and high-latency retrieval. For instance, the Wikitext-103 datastore requires nearly 500 GB of storage even for the GPT2-small model (He et al., 2021). These limitations motivate our MLP Memory, a compact parametric model pretrained to approximate the retrieval function: given a query embedding, it directly outputs a NN-like next token distribution, thereby eliminating both the substantial storage requirements and high-latency retrieval.
@@ 3.1 Architecture
In designing the memory module, we observe from Section 2 that the retriever imitation task processes a single-vector representation without requiring token-mixing operations. Recent studies (Geva et al., 2020) have identified that FFN layers function as key-value memories, suggesting that MLPs play a specialized role in knowledge memorization within LLMs. Based on these insights, we propose pretraining an all-MLP memory that effectively functions as a non-parametric retriever, as illustrated in Figure 4.
@@ 3.2 Training
Loss Function Unlike traditional language modeling with single-label targets, NN distributions capture the diversity of plausible continuations by encoding multiple valid next tokens weighted by their contextual similarity. Our ablation studies in Section 5.4 demonstrate that a hybrid objective combining two complementary losses yields optimal performance. Our approach centers on minimizing the Kullback-Leibler divergence (Van Erven & Harremos, 2014) between MLP Memory’s output distribution and the cached NN distributions:
This encourages the memory module to match the full probability distribution rather than merely predicting the most likely token. To prevent excessive deviation from the underlying corpus distribution, we integrate a complementary Cross-Entropy loss (Zhang & Sabuncu, 2018):
@@ 4 Experimental Setup
Overview We conduct comprehensive experiments to evaluate MLP Memory across four critical dimensions. First, we assess performance on five question-answering benchmarks (5.1) to demonstrate that our approach represents a novel form of parametric memory that surpasses both traditional parametric methods (continued pretraining, LoRA) and non-parametric approaches (RAG). Second, we evaluate on fundamental NLP tasks (5.2) to verify that integrating MLP Memory preserves the base model’s general capabilities. Third, we examine hallucination reduction (5.3) on HaluEval to validate our method’s effectiveness in improving factual accuracy. Finally, we present an ablation study (5.4) to analyze design choices such as loss weighting and layer selection.
@@
Implementation Details We conduct our experiments on 32×A800 80GB GPUs. To demonstrate the generalizability of our approach, we employ two distinct backbone models: Llama-2-7B (Touvron et al., 2023) and Mistral-7B-v0.3 (Jiang et al., 2023). For question-answering benchmarks, we build key-value datastores and non-parametric distributions using both models on preprocessed Wikipedia-2021 (Izacard et al., 2022), and train separate 1B-parameter MLP Memory modules with learning rate 4e-4. For general NLP tasks, we build datastores using Mistral-7B-v0.3 on a heterogeneous corpus following Geng et al. (2024), and train the MLP Memory with learning rate 4e-4. For hallucination evaluation, we directly apply the MLP Memory trained from question-answering experiments. All experiments use a training budget equivalent to the computational cost of training a 7B parameter model for 1 epoch. The training hyperparameter is set to 0.4 across all tasks. The interpolation hyperparameter is tuned on the validation split of each task following Khandelwal et al. (2020), see more details in Appendix D.
@@
Baselines We compare MLP Memory against established methods for improving factual accuracy and knowledge utilization: RAG, which employs BGE (Chen et al., 2024) as the retrieval model and retrieves top-5 documents to ensure comprehensive context coverage. NN-LM (Khandelwal et al., 2020), configured with interpolation parameter and temperature following (Geng et al., 2024). LoRA (Hu et al., 2022), applied to query, key, value, and MLP layers, with rank adjusted to match the parameter count of our MLP Memory modules. Continued Pretraining (CPT), which involves further training of all model parameters on the corresponding corpus.
@@ 5.1 Question Answering Performance
We evaluate MLP Memory on five diverse QA benchmarks: Natural Questions (NQ) (Kwiatkowski et al., 2019), WebQA (Berant et al., 2013), TriviaQA (Joshi et al., 2017), TruthfulQA (Lin et al., 2022), and HotpotQA (Yang et al., 2018), comparing against CPT, LoRA, and RAG. As shown in Table 4, Mistral-7B-v0.3 with MLP Memory achieves an average relative improvement of 12.3% over the baseline across five benchmarks, with particularly striking improvements on NQ (25.20% vs. baseline 20.63%) and WebQA (37.45% vs. baseline 29.28%). While CPT and LoRA suffer significant degradation across all tasks—likely due to catastrophic forgetting during domain-specific training—MLP Memory maintains or improves performance by learning to emulate retrieval behavior without modifying the base model’s parameters. Notably, our approach outperforms RAG despite RAG having access to the same external knowledge at inference time, suggesting that our parametric compression of retrieval patterns captures richer contextual relationships than explicit document retrieval. The consistent gains across both factoid QA (NQ, TriviaQA) and multi-hop reasoning (HotpotQA) demonstrate that MLP Memory effectively bridges the gap between parametric and non-parametric memory systems.
@@ 5.2 General NLP Task Performance
To ensure MLP Memory doesn’t compromise fundamental language understanding, we evaluate on nine standard NLP tasks spanning sentiment classification (SST-2 (Socher et al., 2013), MR (Pang & Lee, 2005a), CR (Hu & Liu, 2004), RT (Pang & Lee, 2005b), HYP (Kiesel et al., 2019)), textual entailment (CB (De Marneffe et al., 2019), RTE (Dagan et al., 2010)), and topic classification (AGNews (Zhang et al., 2015a), Yahoo (Zhang et al., 2015b)). Table 5.1 reveals that MLP Memory achieves comprehensive improvements across all general tasks, achieving the highest average score compared to all baselines. The improvements are particularly pronounced on reasoning-intensive tasks like RTE (64.62% vs. baseline 59.57%) and CB (76.79% vs. baseline 69.64%), suggesting that the retrieval patterns learned by MLP Memory provide useful inductive biases even for tasks that don’t explicitly require factual knowledge retrieval. In contrast, CPT and LoRA show mixed results with improvements on some tasks but degradation on others. The robust performance across this diverse task suite demonstrates that MLP Memory’s external parametric memory complements rather than interferes with the base model’s learned representations.

### v3

@@ Abstract
Modern approaches to enhancing Large Language Models’ factual accuracy and knowledge utilization face a fundamental trade-off: non-parametric retrieval-augmented generation (RAG) provides flexible access to external knowledge but suffers from high inference latency and shallow integration, while parametric fine-tuning methods like LoRA risk catastrophic forgetting and degraded general capabilities. In this work, we propose MLP Memory, a lightweight parametric module that learns to internalize retrieval patterns without explicit document access. By pretraining an MLP to imitate a NN retriever’s behavior on the entire pretraining dataset, we create a differentiable memory component that captures the benefits of retrieval-based knowledge access in a fully parametric form. Our architecture integrates this pretrained MLP Memory with Transformer decoders through simple probability interpolation, yielding 17.5% and 24.1% scaling gains on WikiText-103 and Web datasets, respectively. It further achieves 12.3% relative improvement on five question-answering benchmarks and 5.2 points absolute gain across nine general NLP tasks, while reducing hallucinations by up to 10 points on HaluEval. Moreover, MLP Memory delivers 2.5 faster inference than RAG with superior accuracy. Our findings show that learning retrieval patterns parametrically bridges the gap between efficient inference and effective knowledge access, offering a practical alternative to both RAG and fine-tuning approaches.
@@ 1 Introduction
Decoder-only architectures such as GPT [Brown et al., 2020], LLaMA [Grattafiori et al., 2024], Qwen [Qwen et al., 2025], and DeepSeek [Liu et al., 2024] have achieved remarkable success in various tasks, including open-ended text generation [OpenAI et al., 2024], code completion [Chen et al., 2021], image synthesis [Chen et al., 2020], and multimodal reasoning [Liu et al., 2023]. However, despite their impressive capabilities, these models often struggle with effective knowledge utilization, producing responses that may be fluent but fail to accurately leverage the factual information encoded in their parameters.
@@
Current approaches to enhance knowledge utilization in LLMs face significant trade-offs. Retrieval-augmented generation (RAG) methods [Lewis et al., 2021, Peng et al., 2023, Gao et al., 2022, Izacard et al., 2022] dynamically fetch relevant documents to ground model outputs, providing flexible access to external knowledge sources. However, these non-parametric approaches introduce substantial inference latency through expensive nearest-neighbor searches and longer context from retrieved documents. They also suffer from shallow integration with the base model, as the retrieval component remains isolated from the LLM’s computational graph. Conversely, parametric adaptation methods such as continued pre-training (CPT) and LoRA [Hu et al., 2022] directly modify model weights to incorporate domain-specific knowledge. While computationally efficient at inference time, these approaches risk catastrophic forgetting of previously learned capabilities and often degrade performance on general tasks, requiring careful task-specific tuning that limits their broader applicability. Figure 2 illustrates how our approach differs fundamentally from both non-parametric retrieval methods and parametric adaptation approaches.
@@
In contrast to decoder-only LLMs, neuroscience research reveals a lateralized human brain where language processing is dominated by the left hemisphere while memory formation occurs in the hippocampus [Gazzaniga, 2005b, a, Douglas, 1967]. This insight has inspired memory-augmented models in machine learning. Early approaches like Memory Networks [Weston et al., 2015] enabled read/write operations on external memory, while Sparse Access Memory introduced differentiable memory access schemes. However, these were task-specific with limited general applicability. In the LLM era, methods such as Memory Transformers [Burtsev et al., 2021] incorporate trainable memory tokens for global context, while AutoCompressors [Chevalier et al., 2023] compress long contexts into summary vectors. Nevertheless, these memory tokens primarily function as working memory supplements for context extension rather than long-term memory capable of retaining information from the entire training corpus.
@@
In this work, we propose an external memory for LLM that is pretrained to mimic a retriever on the entire pretraining dataset. Specifically, following the RAG setting in NN-LM [Khandelwal et al., 2020], this memory learns to map the LLM hidden state at a certain step to a vocabulary distribution matching the output of the NN retriever. During inference, the LLM’s native output is interpolated with the retriever-pretrained output from the external memory. Our resulting architecture, illustrated in Figure 4, consists of a transformer decoder and an external MLP memory, each pretrained separately with different pretraining tasks. For our pretrained external memory, we aim to achieve the following features simultaneously:
@@
Scalable, general knowledge memorization, covering the whole training set. MLP memory is trained on the entire pretraining dataset, not limited to the context level. It supports general knowledge retention and demonstrates stronger scaling behavior than decoder-only models.
5)
Long-term memory. While existing memory tokens serve primarily as working memory by storing local context for immediate use, our MLP memory functions as a long-term repository of generalizable knowledge acquired during the pretraining phase.
Experimental results demonstrate that MLP Memory significantly outperforms existing approaches across multiple dimensions. Our architecture exhibits steeper power-law scaling with model size, achieving 17.5% and 24.1% improvement on WikiText-103 and Web datasets compared to decoder-only architectures, while continuing to benefit from additional training without overfitting. On downstream tasks, it achieves average relative improvements of 12.3% (Mistral-7B) and 7.8% (Llama2-7B) on five QA benchmarks, with WebQA showing exceptional gains (37.45% vs. 29.28% baseline). On nine general NLP tasks, it delivers a 5.2 points absolute improvement. MLP Memory also substantially reduces hallucinations on HaluEval, with accuracy improvements of 9.68, 10.08, and 2.14 points on dialogue, QA, and summarization tasks respectively. Most notably, it achieves 2.5 faster time-to-first-token than RAG and 5.6 faster than NN-LM, while maintaining constant inference speed regardless of corpus size, unlike retrieval methods whose latency scales with data size. Figure 1 illustrates MLP Memory’s performance gains and inference efficiency over baselines and Figure 3 demonstrates a case where MLP Memory correctly answers factual questions while RAG fails despite retrieving correct information. These results confirm that parametric compression of retrieval patterns offers a more efficient and effective alternative to explicit retrieval.
@@ 2 Preliminary: -nearest neighbors language model
The NN-LM [Khandelwal et al., 2020] augments a pre-trained LM by interpolating its parametric distribution with a non-parametric distribution from nearest neighbor retrieval. Given context , and denotes the next token. The next-token probability is:
@@
While NN-LM improves predictions through explicit memory, it suffers from substantial storage requirements and high-latency retrieval. For instance, the Wikitext-103 datastore requires nearly 500 GB of storage even for the GPT2-small model [He et al., 2021]. These limitations motivate our MLP Memory, a compact parametric model pretrained to approximate the retrieval function: given a query embedding, it directly outputs a NN-like next token distribution, thereby eliminating both the substantial storage requirements and high-latency retrieval.
@@ 3.1 Architecture
In designing the memory module, we observe from Section 2 that the retriever imitation task processes a single-vector representation without requiring token-mixing operations. Recent studies [Geva et al., 2020] have identified that FFN layers function as key-value memories, suggesting that MLPs play a specialized role in knowledge memorization within LLMs. Based on these insights, we propose pretraining an all-MLP memory that effectively functions as a non-parametric retriever, as illustrated in Figure 4.
@@ 3.2 Training
Loss Function Unlike traditional language modeling with single-label targets, NN distributions capture the diversity of plausible continuations by encoding multiple valid next tokens weighted by their contextual similarity. Our ablation studies in Section 5.5 demonstrate that a hybrid objective combining two complementary losses yields optimal performance. Our approach centers on minimizing the Kullback-Leibler divergence [Van Erven and Harremos, 2014] between MLP Memory’s output distribution and the cached NN distributions:
This encourages the memory module to match the full probability distribution rather than merely predicting the most likely token. To prevent excessive deviation from the underlying corpus distribution, we integrate a complementary Cross-Entropy loss [Zhang and Sabuncu, 2018]:
@@ 4 Experimental Setup
Overview We conduct comprehensive experiments to evaluate MLP Memory across five critical dimensions. First, we analyze the scaling behavior (5.1) of models augmented with MLP memory through power-law fitting. Second, we assess performance on five question-answering benchmarks (5.2) to demonstrate that our approach represents a novel form of parametric memory that surpasses both traditional parametric methods (continued pretraining, LoRA) and non-parametric approaches (RAG). Third, we evaluate on fundamental NLP tasks (5.3) to verify that integrating MLP Memory preserves the base model’s general capabilities. Fourth, we examine hallucination reduction (5.4) on HaluEval to validate our method’s effectiveness in improving factual accuracy. Finally, we present an ablation study (5.5) to analyze design choices such as loss weighting and layer selection.
@@
Implementation Details We conduct our experiments on 32×A800 80GB GPUs. To demonstrate the generalizability of our approach, we employ two distinct backbone models: Llama-2-7B [Touvron et al., 2023] and Mistral-7B-v0.3 [Jiang et al., 2023]. For scaling law, please refer to Section 5.1 for detailed settings. For question-answering benchmarks, we build key-value datastores and non-parametric distributions using both models on preprocessed Wikipedia-2021 [Izacard et al., 2022], and train separate 1B-parameter MLP Memory modules with learning rate 4e-4. For general NLP tasks, we build datastores using Mistral-7B-v0.3 on a heterogeneous corpus following Geng et al. [2024], and train the MLP Memory with learning rate 4e-4. For hallucination evaluation, we directly apply the MLP Memory trained from question-answering experiments. All experiments use a training budget equivalent to the computational cost of training a 7B parameter model for 1 epoch. The training hyperparameter is set to 0.4 across all tasks. The interpolation hyperparameter is tuned on the validation split of each task following Khandelwal et al. [2020], see more details in Appendix B.
@@
Baselines We compare MLP Memory against established methods for improving factual accuracy and knowledge utilization: RAG, which employs BGE [Chen et al., 2024] as the retrieval model and retrieves top-5 documents to ensure comprehensive context coverage. NN-LM [Khandelwal et al., 2020], configured with interpolation parameter and temperature following [Geng et al., 2024]. LoRA [Hu et al., 2022], applied to query, key, value, and MLP layers, with rank adjusted to match the parameter count of our MLP Memory modules. Continued Pretraining (CPT), which involves further training of all model parameters on the corresponding corpus.
5 Experimental Results
5.1 Scaling law
Setup
We conduct scaling law experiments using standard decoder-only models and our overall model architecture. As baselines, we use four GPT-2 Radford et al. [2019] variants with increasing parameter counts: GPT2-small (124M), GPT2-medium (345M), GPT2-large (774M), and GPT2-xl (1.5B). For MLP Memory, we define three configurations: small (124M), medium (335M), and large (774M) that align with the scaling trend of standard architectures. The MLP Memory module is externally integrated with a matching-sized GPT-2 variant, resulting in total parameter counts of approximately 248M, 710M, and 1.5B for the small, medium, and large configurations, respectively. All models are trained on two datasets: WikiText-103 Merity et al. [2016] (around 100M tokens) and a mixed Web dataset (around 600M tokens). Following Shi et al. [2022a], our Web dataset combines diverse knowledge sources relevant to common NLP tasks, including WikiText-103, Amazon Reviews He and McAuley [2016], CC-NEWS Hamborg et al. [2017], and IMDB Maas et al. [2011].
Scaling law with model parameters
Following Kaplan et al. [2020], we model perplexity scaling as . Under fixed compute, we compare our architecture to continued GPT-2 training on WikiText-103 and Web datasets in terms of test perplexity scaling with model size . Results in Figure 5 show our architecture demonstrates a steeper scaling curve than the decoder-only model, indicating improved scaling efficiency. The power-law scaling laws on WikiText-103 can be expressed as:
Scaling law with training compute
@@ 5.2 Question Answering Performance
We evaluate MLP Memory on five diverse QA benchmarks: Natural Questions (NQ) [Kwiatkowski et al., 2019], WebQA [Berant et al., 2013], TriviaQA [Joshi et al., 2017], TruthfulQA [Lin et al., 2022], and HotpotQA [Yang et al., 2018], comparing against CPT, LoRA, and RAG. As shown in Table 1, Mistral-7B-v0.3 with MLP Memory achieves an average relative improvement of 12.3% over the baseline across five benchmarks, with particularly striking improvements on NQ (25.20% vs. baseline 20.63%) and WebQA (37.45% vs. baseline 29.28%). While CPT and LoRA suffer significant degradation across all tasks—likely due to catastrophic forgetting during domain-specific training—MLP Memory maintains or improves performance by learning to emulate retrieval behavior without modifying the base model’s parameters. Notably, our approach outperforms RAG despite RAG having access to the same external knowledge at inference time, suggesting that our parametric compression of retrieval patterns captures richer contextual relationships than explicit document retrieval. The consistent gains across both factoid QA (NQ, TriviaQA) and multi-hop reasoning (HotpotQA) demonstrate that MLP Memory effectively bridges the gap between parametric and non-parametric memory systems.
@@ 5.3 General NLP Task Performance
To ensure MLP Memory doesn’t compromise fundamental language understanding, we evaluate on nine standard NLP tasks spanning sentiment classification (SST-2 [Socher et al., 2013], MR [Pang and Lee, 2005a], CR [Hu and Liu, 2004], RT [Pang and Lee, 2005b], HYP [Kiesel et al., 2019]), textual entailment (CB [De Marneffe et al., 2019], RTE [Dagan et al., 2010]), and topic classification (AGNews [Zhang et al., 2015a], Yahoo [Zhang et al., 2015b]). Table 2 reveals that MLP Memory achieves comprehensive improvements across all general tasks, achieving the highest average score compared to all baselines. The improvements are particularly pronounced on reasoning-intensive tasks like RTE (64.62% vs. baseline 59.57%) and CB (76.79% vs. baseline 69.64%), suggesting that the retrieval patterns learned by MLP Memory provide useful inductive biases even for tasks that don’t explicitly require factual knowledge retrieval. In contrast, CPT and LoRA show mixed results with improvements on some tasks but degradation on others. The robust performance across this diverse task suite demonstrates that MLP Memory’s external parametric memory complements rather than interferes with the base model’s learned representations.